Circular No. 7011

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dx IN UGC 11149 M. Modjaz, T. Shefler, E. Halderson, J. Y. King, W. D. Li, R. R.

Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of a supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998dx was found and confirmed on unfiltered images taken on Sept. 10.2 (mag ∼ 18.3) and 12.2 UT (mag ∼ 17.9), respectively. The object is located at α = 18^{h}11^{m}11^{s}.89, δ = +49^{o}51′40′′.7 (equinox 2000.0), which is ∼ 20′′.9 east and 12′′.2 south of the nucleus of UGC 11149. A KAIT image obtained on Sept. 7.2 showed the new star at mag 18.8, and an image taken on Sept. 3.2 already showed a hint of the supernova. There is nothing at the position of SN 1998dx (limiting mag ∼ 19.0) in a KAIT image obtained on July 30.3.

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1998dx taken by P. Berlind with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) on Sept. 14.2 UT reveals that it is a type-Ia supernova near maximum light. No direct host-galaxy recession velocity information is available, but the prominent Si II (rest 635.5 nm) absorption feature appears at 645 nm. Assuming a photospheric expansion velocity typical of maximum-light type-Ia supernovae, such that the feature would appear at 615 nm in the host-galaxy frame, the observed wavelength would imply a host-galaxy recession velocity of 15 000 km/s.

SUPERNOVA 1998dt IN NGC 945 Jha et al. also write: “A spectrum of SN 1998dt (cf. IAUC 7006)

taken by M. Calkins on Sept. 12.4 UT (equipment as above) exhibits a blue continuum and strong absorption lines of He I and Fe II, indicating that the supernova is of type Ib near maximum light. The host galaxy recession velocity is 4600 km/s, measured from narrow emission lines of Hα.”

SUPERNOVA 1998dh IN NGC 7541 CCD V magnitude estimates (cf. IAUC 6989): Aug. 15.957 UT, 14.2

(D. Hanžl, Brno, Czech Republic); 25.854, 14.65 (Hanžl); 27.974, 15.0 (D. Rodriguez, Madrid, Spain); Sept. 9.80, 15.70 ± 0.05 (L. Kiss, Konkoly Observatory, Piszkéstető). Unfiltered CCD magnitude estimates (cf. IAUC 6999): July 29.87, 14.3 (S. Moretti and S. Tomaselli, Forlí, Italy); Aug. 30.94, 15.0 (L. Boschini, Miragolo, Bergamo, Italy).

1998 September 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green